OpenStudy (anonymous):

Need help on this one please. Solve for y using the multiplication principle: y/5= -6?

OpenStudy (anonymous):

(y)/(5)=-6 Multiply each term in the equation by 5. (y)/(5)*5=-6*5 Cancel the common factor of 5 in the denominator of the first term (y)/(5) and the second term 5. (y)/(<X>5<x>)*<X>5<x>=-6*5 Reduce the expression by removing the common factor of 5 in the denominator of the first term (y)/(5) and the second term 5. y*1=-6*5 Multiply y by 1 to get y. y=-6*5 Multiply -6 by 5 to get -30. y=-30

OpenStudy (anonymous):

Thank you Some. I am trying to do this in my own way but you make it look easier via your way :0)